# ASVD: Activation-aware Singular Value Decomposition for Compressing Large Language Models

### v1

@@ 3.3 Activation-aware Singular Value Decomposition
This scaling step is crucial as it modifies the weight matrix to reflect the varying significance of different input channels, setting the stage for a more effective and targeted decomposition process.
@@
This step is pivotal as it ensures that the most significant aspects of the scaled weight matrix are retained, while less critical information, which contributes minimally to the model’s output, is discarded.
@@
Our ASVD, with its focus on channel-wise activation outliers, strives to mitigate these outliers by appropriately scaling the corresponding weights. In this way, the activation variation can be mitigated as illustrated in Fig. 2. Note that the diagonal nature of $\mathbf{S}$ simplifies its inversion process, making $\mathbf{S}^{-1}$ computationally efficient. The subsequent challenge is to devise a method or criteria for accurately determining $\mathbf{S}$ so that it effectively aligns with the activation patterns observed in the LLM, ensuring that the scaling process optimally addresses the outlier activations.
Scaling Matrix $\mathbf{S}$ Exploration. Each diagonal element in the scaling matrix $\mathbf{S}$, particularly $\mathbf{S}_{ii}$ is pivotal in modulating the input activations as $(\mathbf{W}\mathbf{S})\mathbf{X}$. Each entry of $\mathbf{S}_{ii}$ specifically scales the influence of the $i$-th input channel. This scaling is instrumental in adjusting how each activation channel impacts the weight matrix during the decomposition process. The significance of $\mathbf{S}_{ii}$ is visualized in Fig.2 (bottom line), where its role in modulating channel influence is depicted. To effectively determine the values of $\mathbf{S}_{ii}$, various methodologies are explored. Each of these methods focuses on different characteristics of input activation, aiming to optimize $\mathbf{S}$ so that it accurately reflects the importance and influence of each input channel. This exploration is crucial for ensuring that the scaling aligns with the activation patterns observed in the LLM, thereby enhancing the effectiveness of the ASVD process.
Absolute Mean Value of Input Activation: This method calculates $\mathbf{S}_{ii}$ by calculating the absolute mean value of the activations in the $i$-th channel. Mathematically, if $\mathbf{X}_{i,:}$ represents the input activation values of the $i$-th channel, then $\mathbf{S}_{ii}$ is computed as:
$\mathbf{S}_{ii}=(\frac{1}{n}\sum_{j=1}^{n}|\mathbf{X}_{ij}|)^{\alpha},$ (8)
@@
where $n$ is the total number of activations for the $i$-th channel and hyper-parameter $\alpha$ provides flexibility to adjust the level of activation sensitivity incorporated into the scaling. This method focuses on the average magnitude of activation in each channel, capturing the general intensity of activation signals irrespective of their positive or negative nature.
Absolute Maximum Value of Input Activation: Diverging from the mean-based approach, this method zeroes in on the peak activation within each channel. Under this approach, $\mathbf{S}_{ii}$ is determined by the highest absolute value observed in the activations of the $i$-th channel. The formal expression for this calculation is:
$\mathbf{S}_{ii}=(\max_{j}(|\mathbf{X}_{ij}|))^{\alpha}.$ (9)
This method places emphasis on the most significant, or peak, activation in each channel, positing it as a key indicator of the channel’s overall impact and importance. By focusing on the peak values, this approach underscores the channels with the most pronounced activation responses, considering them as crucial contributors to the network’s functionality.
@@ 3.4 Sensitivity-based Truncation Rank Searching
In addressing the second challenge - the variations of singular values among different layers of LLMs - we introduce the Sensitivity-based Truncation Rank Searching (STRS) method. This challenge arises from the fact that different layers in LLMs exhibit varying degrees of sensitivity to information compression, which is reflected in the distribution of their singular values. STRS aims to analyze and leverage this layer-specific sensitivity, particularly concerning how truncation of singular values impacts prediction accuracy.
Similar to any post-training compression methods [Dettmers et al., 2022, Frantar et al., 2022, Frantar & Alistarh, 2023], we first collect a small but representative calibration dataset. This dataset is instrumental in evaluating the layers’ responsiveness to different truncation ranks. For the calibration set, we select a sample of sequences from a source such as Wikitext [Merity et al., 2016]. In the realm of NLP, perplexity is a key metric for assessing how effectively a language model predicts a sequence of tokens [Brown et al., 2020]. Within the scope of our analysis, “sensitivity” refers to the reduction in perplexity observed in the calibration dataset following the decomposition of the model.
The core of the sensitivity evaluation process involves an in-depth exploration of how the neural network reacts to varying levels of truncation. We define a set of potential truncation ratios, denoted as $R$, which include values like $\{0.1,0.2,\cdots,0.9\}$. These ratios determine the fraction of the rank $k$ preserved during the SVD truncation for a weight matrix with dimensions $m\times n$. The truncation ratio $r$ is mathematically defined as:
$r=\frac{km+kn}{mn},r\in R.$ (10)
For each linear layer in the LLM, we iterate through these candidate ratios. At each ratio, truncated SVD is applied to the layer’s weight matrix, temporarily replacing the original layer in the model with this truncated version. The model’s perplexity is then evaluated on the calibration dataset with these adjustments. This iterative process allows us to determine the truncation ratio that optimally balances model compression with minimal impact on prediction accuracy.
@@
This detailed exploration of sensitivity across various truncation levels provides essential insights into each layer’s performance dynamics, informing the optimization and decision-making processes in model compression. As illustrated in Fig. 3, there are noticeable variations in sensitivity among the different layers of the network. Three key observations emerge from this analysis: (i) Inversely Proportional Relationship: We observe that perplexity is inversely proportional to the parameter ratio. Specifically, lower parameter ratios tend to result in higher perplexity scores. This finding highlights the critical trade-off between model compression and reasoning performance, signaling the importance of careful parameter selection to maintain a balance. (ii) Higher Sensitivity in MLP Layers: Layers within Multi-Layer Perceptrons (MLP) [Vaswani et al., 2017] demonstrate higher sensitivity compared to those in feed forward modules. This distinction is crucial in guiding the compression strategy, indicating where more cautious truncation is necessary. (iii) Variable Sensitivity Among Layers: Some layers exhibit relatively lower sensitivity, indicating potential for more aggressive compression without significant performance degradation. This variation suggests the presence of redundancy within LLMs, presenting opportunities for more efficient compression approaches. Targeting these less-sensitive layers could lead to models that are both compact and computationally efficient.
Post sensitivity evaluation, we compile a list of tuples (layer, truncation rank, sensitivity) and sort it according to sensitivity. To determine an optimal rank configuration that aligns with these insights, we propose a binary search algorithm. The details of this algorithm are outlined in Algorithm 1, designed to navigate the complexity of selecting the most effective truncation ranks for each layer, based on their sensitivity and contribution to the overall model performance. More detailed explanations about this algorithm can be found in the Appendix.
3.5 Absorbing Singular Values
param ratio weight quant absorbed by UV absorbed by U absorbed by V
0.9 INT6 37.58 39.67 40.62
0.85 INT6 60.44 64.19 61.02
@@ 4.1 Hyper-parameters Exploration
In our study, we initiated an exploration of hyper-parameters in ASVD, focusing on the activation channel significance metric and the control factor $\alpha$. This exploration were conducted on OPT-125m, a relatively small network that facilitates rapid evaluation.
4.2 Evaluation of ASVD on LLaMA Models
We conducted a comprehensive evaluation of Adaptive Singular Value Decomposition (ASVD) on two LLaMA models: LLaMA-7b and LLaMA-2-7b, each comprising 7 billion parameters. For each model, we selected 32 calibration datasets, each containing 2048 tokens, from the Wikitext dataset to assess the layer-wise sensitivity.
$\alpha$ 0.1 0.25 0.5 1 2
SVD+STRS 103.39
ASVD abs mean 47.54 37.12 36.89 41.53 43.81
ASVD abs max 52.63 47.17 40.14 41.94 52.55
@@ 4.2 Evaluation of ASVD on LLaMA Models
Table 3: Performance under various compression scenarios. Param ratio indicates the proportion of parameters remaining after decomposition. MMLU results are 0-shot.
method param ratio MMLU wiki ptb
LLaMA-7b
original 1 30.76% 5.68 29.63
SVD 0.95 22.98% 2800.94 5458.67
SVD+STRS 0.95 23.92% 136.05 183.92
SVD+STRS 0.9 23.54% 698.66 262.03
ASVD 0.95 30.26% 5.78 32.64
ASVD 0.9 29.67% 6.09 37.80
ASVD 0.85 29.70% 6.80 52.11
ASVD 0.8 27.85% 8.89 88.09
ASVD 0.75 24.94% 14.51 212.80
LLaMA-2-7b
original 1 34.86% 5.47 20.82
SVD 0.95 - nan nan
SVD+STRS 0.95 24.78% 46.79 363.37
SVD+STRS 0.9 24.31% 114.45 27660.43
ASVD 0.95 33.24% 5.64 23.98
ASVD 0.9 32.58% 5.93 32.63
ASVD 0.85 31.57% 6.74 59.84
ASVD 0.8 28.15% 8.91 114.70
ASVD 0.75 25.97% 18.97 432.57
LLaMA-2-13b
original 1 40.16% 4.88 29.21
SVD 0.95 - nan nan
SVD+STRS 0.95 24.86% 167.63 567.02
SVD+STRS 0.9 - nan nan
ASVD 0.95 39.52% 4.94 31.93
ASVD 0.9 40.04% 5.12 34.03
ASVD 0.85 37.95% 5.54 39.32
ASVD 0.8 34.63% 6.53 59.68
ASVD 0.75 28.59% 8.71 110.10
Our methodology involved setting varying thresholds binary searching process, enabling us to observe the impact of different compression levels on model performance. This approach resulted in a range of compressed networks, each characterized by a unique compression ratio.
@@
We evaluated the performance of these compressed networks using perplexity as the primary metric, focusing on two datasets: Wikitext-2 and the Penn Treebank (PTB). The objective was to understand the relationship between compression ratio and model performance. The results, illustrated in Figure 4, reveal several key insights: (1) As the parameter ratio decreases, there is a corresponding increase in perplexity. As more parameters are compressed, the more error is introduced by SVD. This reduces representational capacity of the network, thereby resulting in higher perplexity scores. (2) A plateau region is observed when the parameter ratio exceeds 0.9. In this range, ASVD predominantly decompresses the less sensitive layers, resulting in minimal impact on prediction accuracy. (3) Below a parameter ratio of 0.85, there is a rapid increase in perplexity, indicating that the more sensitive layers are being decompressed to a lower truncation rank, adversely affecting the model’s performance.
We present a detailed analysis of the performance of compressed networks at various parameter ratios, as summarized in Table 3. This analysis aims to elucidate the impact of different compression techniques on network efficiency and accuracy.
Table 3 displays the performance metrics for two LLaMA models, LLaMA-7b and LLaMA-2-7b, under several compression scenarios. These metrics include Massive Multitask Language Understanding (MMLU) zero-shot evaluation, perplexity on the Wikitext dataset (wiki), and perplexity on the Penn Treebank dataset (ptb).
@@
Our observations indicate significant variations in performance as a function of the parameter ratio and the compression method employed. Specifically, the table highlights the performance of each model when using standard Singular Value Decomposition (SVD), SVD with binary search for truncation ranks (SVD*), and ASVD at different parameter ratios ranging from 0.75 to 0.95.
We also compare ASVD with FWSVD [Hsu et al., 2022], which uses Fisher information to weigh the importance of parameters affecting the model prediction. Note that FWSVD is training-required. As shown in Table 4, our method can outperform FWSVD comprehensively.
param ratio 0.95 0.9 0.85 0.8
LLaMA-7b
FWSVD+STRS wiki 5.86 6.32 7.48 10.70
ASVD 5.78 6.09 6.80 8.89
FWSVD+STRS ptb 34.33 38.05 58.75 125.80
ASVD 32.64 37.80 52.11 88.09
LLaMA-2-7b
FWSVD+STRS wiki 5.59 6.12 8.01 13.07
ASVD 5.64 5.93 6.74 8.91
FWSVD+STRS ptb 25.06 36.58 105.53 222.03
ASVD 23.98 32.63 59.84 114.70
@@ 4.3 Integrating ASVD with Quantization
param ratio dtype LLaMA-2-7b LLaMA-2-13b
wiki ptb wiki ptb
1 FP16 5.47 20.82 4.88 29.15
1 INT8 5.48 20.82 4.88 29.12
1 INT6 5.52 20.97 4.90 29.03
1 NF4 5.65 22.70 4.98 30.31
0.95 FP16 5.64 23.98 4.94 31.93
0.95 INT8 5.64 23.95 4.95 31.67
0.95 INT6 5.70 25.10 4.97 31.97
0.95 NF4 5.83 35.91 5.08 33.89
0.9 FP16 5.93 32.63 5.12 34.03
0.9 INT8 5.94 32.19 5.11 33.64
0.9 INT6 6.02 32.83 5.16 33.66
0.9 NF4 6.20 40.82 5.31 34.93
0.85 FP16 6.74 59.84 5.54 39.32
0.85 INT8 6.73 63.76 5.56 40.02
0.85 INT6 6.90 66.88 5.62 38.15
0.85 NF4 7.43 427.59 5.90 44.49
This section explores the integration of ASVD with quantization techniques to compress large language models. While our primary focus is on ASVD, we investigate the synergy between ASVD and simple quantization methods, Round-To-Nearest (RTN) and 4-bit NormalFloat (NF4) [Dettmers et al., 2023], as a preliminary step. Future work could extend this investigation to more advanced quantization methods.
Our experimental framework involves two stages. Firstly, we apply ASVD to decompose the network. Subsequently, we quantize the decomposed weights. We employ per-channel asymmetric quantization, targeting two quantization levels: 8-bit and 6-bit. We employ NF4 for quantizing to 4-bit. The impact of these quantization levels on network performance is then assessed.

### v3

@@ 3.3 ASVD: Activation-aware Singular Value Decomposition
This scaling step is crucial as it modifies the weight matrix to reflect the varying significance of different input channels, making the decomposition more effective and targeted.
@@
This step is pivotal as it ensures that the most significant aspects of the scaled weight matrix are retained. While less critical information, which contributes minimally to the model’s output, is discarded.
@@
ASVD focus on channel-wise activation outliers, striving to mitigate these outliers by appropriately scaling the corresponding weights. In this way, the activation variation can be mitigated as illustrated in Fig. 2. Note that the diagonal nature of $\mathbf{S}$ simplifies its inversion process, making $\mathbf{S}^{-1}$ computationally efficient. The subsequent challenge is to devise a method or criteria for accurately determining $\mathbf{S}$ so that it effectively aligns with the activation patterns observed in the LLM, ensuring that the scaling process optimally addresses the outlier activations.
$S_{ii}$ is pivotal in moduling the input channel as: $(WS)_{:,i}=W_{:,i}S_{ii}$
@@
Setting Scaling Matrix $\mathbf{S}$. Each diagonal element in the scaling matrix $\mathbf{S}$, particularly $\mathbf{S}_{ii}$ is pivotal in moduling the input channel as: $(\mathbf{WS})_{:,i}=\mathbf{W}_{:,i}\mathbf{S}_{ii}$. Each entry of $\mathbf{S}_{ii}$ specifically scales the influence of the $i$-th input channel. This scaling is instrumental in adjusting how each activation channel impacts the weight matrix during the decomposition process. The significance of $\mathbf{S}_{ii}$ is visualized in Fig.2 (bottom line), where its role in modulating channel influence is depicted. We use absolute mean value of input activation to set $\mathbf{S}_{ii}$. We calculates $\mathbf{S}_{ii}$ by calculating the absolute mean value of the activations in the $i$-th channel. Mathematically, if $\mathbf{X}_{i,:}$ represents the input activation values of the $i$-th channel, then $\mathbf{S}_{ii}$ is computed as: $\mathbf{S}_{ii}=(\frac{1}{n}\sum_{j=1}^{n}|\mathbf{X}_{ij}|)^{\alpha}$, where $n$ is the total number of activations for the $i$-th channel and hyper-parameter $\alpha$ provides flexibility to adjust the level of activation sensitivity incorporated into the scaling. This method focuses on the average magnitude of activation in each channel, capturing the general intensity of activation signals irrespective of their positive or negative nature.
3.4 KV Cache Compression
Figure 3: A demonstration of how ASVD reduces the memory cost of the K cache. (Left) With long text lengths $L$, the memory required for storing the K cache in a $N$-dimensional space becomes substantial. (Right) ASVD decomposes the key projection weight matrix $\mathbf{W}$ into two low-rank matrices $\mathbf{U}$ and $\mathbf{V}$ (see Sec. 3.3). This low-rank structure allows the K representation to be stored in a reduced $r$-dimensional space, where $r\ll N$. Consequently, we only need to save the intermediate K in the $r$ dimension instead of $N$ dimension, saving the K cache $\frac{N}{r}$ times. Note that saving V cache is the same, and when content length $L$ becomes really large (e.g., 1M tokens) or with larger batch size, the KV cache becomes a significant factor in memory cost.
LLM inference with large context lengths can be incredibly resource-intensive, requiring high-end GPUs and, for the largest models, costly multi-GPU setups. Analysis of generative inference with LLMs reveals that, for relatively small batch sizes, the computation is primarily memory-bound [Hooper et al., 2024, Liu et al., 2024]. Given the growing gap between computational speeds and memory speeds, this issue is expected to worsen over time, making it crucial to address the memory bottleneck. Further analysis indicates that the memory bottleneck is strongly correlated with context size. For long sequence lengths, the main contributor to memory consumption is the KV cache storing, so minimizing the KV cache can reduce both memory consumption and bandwidth requirements [Yuan et al., 2024].
As we discussed in Sec.3.3, ASVD decomposes the key and value projection weight matrix $\mathbf{W}\in\mathbb{R}^{N\times N}$ into two low-rank matrices, $\mathbf{U}\in\mathbb{R}^{N\times r}$ and $\mathbf{V}\in\mathbb{R}^{N\times r}$, in a training-free manner, where $N$ is the dimension of K/V embedding space. As shown in Fig.3, replacing the high-rank matrix with two low-rank matrices via ASVD allows us to obtain intermediate activations in low-rank form. These intermediate activations can be stored as a replacement for the original KV cache. In other words, the original KV cache requires storing two $L\times N$ matrices. With ASVD, the new KV cache only needs to store two $L\times r$ matrices. In summary, ASVD can compress the KV cache $\frac{N}{r}$ times. This significant reduction in memory usage for the KV cache enables larger batch sizes or longer sequence lengths, which are critical for real-world applications.
@@ 3.5 Sensitivity-based Truncation Rank Searching
In addressing the second challenge, the variations of singular values among different layers of LLMs, we introduce the Sensitivity-based Truncation Rank Searching (STRS) method. This challenge arises from the fact that different layers in LLMs exhibit varying degrees of sensitivity to information compression, which is reflected in the distribution of their singular values. STRS evaluates the layer sensitivity and decides the best truncation of singular values. In the realm of NLP, perplexity is a key metric for assessing how effectively a language model predicts a sequence of tokens [Brown et al., 2020]. Therefore, we use the reduction in perplexity on the calibration dataset to evaluate the sensitivity of each layer. Similar to post-training compression methods [Dettmers et al., 2022, Frantar et al., 2022, Frantar & Alistarh, 2023], we collect a small number of input token sequences as calibration dataset.
The core of the sensitivity evaluation process involves an in-depth exploration of how the neural network reacts to varying levels of truncation. We define a set of potential truncation ratios, denoted as $R=\{0.1,0.2,\cdots,0.9\}$. These ratios $r=\frac{km+kn}{mn}$ determine the fraction of the rank $k$ preserved during the SVD truncation for a weight matrix with dimensions $m\times n$. For each linear layer in the LLM, we iterate through these candidate ratios. At each ratio, truncated SVD is applied to the layer’s weight matrix, temporarily replacing the original layer in the model with this decomposed version. The model’s perplexity is then evaluated on the calibration dataset.
This detailed exploration of sensitivity across various truncation levels provides essential insights into each layer’s performance dynamics, informing the optimization and decision-making processes in model compression. As illustrated in Fig. 4, there are noticeable variations in sensitivity among the different layers. Three key observations emerge from this analysis: 1. Inversely Proportional Relationship: lower parameter ratios tend to result in higher perplexity scores. 2. Higher Sensitivity in MLP Layers: MLP layers demonstrate higher sensitivity, indicating where more cautious truncation is necessary. 3. Variable Sensitivity Among Layers: Some layers exhibit relatively lower sensitivity, indicating potential for more aggressive compression.
Assuming the affects of layers are independent, we should set the truncation rank of each layer to minimize the total affect to perplexity under the constraint of parameter size. We propose a binary search algorithm to search for the best truncation rank. Detailed explanations of algorithm can be found in the Appendix.
@@ 4.1 Hyper-parameters Exploration
In our study, we initiate an exploration of hyper-parameters in ASVD, focusing on the activation channel significance metric and the control factor $\alpha$. This exploration is conducted on OPT-125m, a relatively small network that facilitates rapid evaluation. We rigorously explored the control factor $\alpha$, and two methods for quantifying activation significance: Absolute Mean Value of Input Activation and Absolute Maximum Value of Input Activation. More details can be found in the Appendix. In summary, based on our explorations, we chose $\alpha=0.5$ and the Absolute Mean Value method for setting the scaling matrix $S$ in the ASVD process in the following experiments.
4.2 Parameters Compression
We conducted a comprehensive evaluation of Adaptive Singular Value Decomposition (ASVD) on two LLaMA models: LLaMA-7b and LLaMA-2-7b [Touvron et al., 2023a, b], each comprising 7 billion parameters. For each model, we selected a calibration set with 32 samples, and each sample contains 2048 tokens, from the Wikitext dataset to assess the layer-wise sensitivity.
@@ 4.2 Parameters Compression
Table 1: Performance under various compression scenarios. Param ratio indicates the proportion of parameters remaining after decomposition. MMLU results are 0-shot. SVD* means SVD using Sensitivity-based Truncation Rank Searching.
LLaMA-7b LLaMA-2-7b LLaMA-2-13b
method param ratio MMLU wiki ptb MMLU wiki ptb MMLU wiki ptb
original 1 30.76% 5.68 29.63 34.86% 5.47 20.82 40.16% 4.88 29.21
SVD 0.95 22.98% 2800 5458 - nan nan - nan nan
SVD* 0.95 23.92% 136.05 183.92 24.78% 46.79 363.37 24.86% 167.63 567.02
SVD* 0.9 23.54% 698.66 262.03 24.31% 114.45 27660 - nan nan
ASVD 0.95 30.26% 5.78 32.64 33.24% 5.64 23.98 39.52% 4.94 31.93
ASVD 0.9 29.67% 6.09 37.80 32.58% 5.93 32.63 40.04% 5.12 34.03
ASVD 0.85 29.70% 6.80 52.11 31.57% 6.74 59.84 37.95% 5.54 39.32
ASVD 0.8 27.85% 8.89 88.09 28.15% 8.91 114.70 34.63% 6.53 59.68
ASVD 0.75 24.94% 14.51 212.80 25.97% 18.97 432.57 28.59% 8.71 110.10
@@
ASVD involves setting varying thresholds binary searching process, enabling us to observe the impact of different compression levels on model performance. This approach resulted in a range of compressed networks, each characterized by a unique compression ratio. We evaluated the performance of these compressed networks using perplexity as the primary metric, focusing on two datasets: Wikitext-2 and the Penn Treebank (PTB). The objective was to understand the relationship between compression ratio and model performance. The results, illustrated in Fig.5, reveal several key insights: (1) As the parameter ratio decreases, there is a corresponding increase in perplexity. As more parameters are compressed, the more error is introduced by SVD. This reduces representational capacity of the network, thereby resulting in higher perplexity scores. (2) A plateau region is observed when the parameter ratio exceeds 0.9. In this range, ASVD predominantly decompresses the less sensitive layers, resulting in minimal impact on prediction accuracy. (3) Below a parameter ratio of 0.85, there is a rapid increase in perplexity, indicating that the more sensitive layers are being decompressed to a lower truncation rank, adversely affecting the model’s performance.
@@
We also present a detailed analysis of the performance of compressed networks at various parameter ratios. This analysis aims to elucidate the impact of different compression techniques on network efficiency and accuracy. Table 1 displays the performance metrics for two LLaMA models, LLaMA-7b and LLaMA-2-7b, under several compression scenarios. These metrics include Massive Multitask Language Understanding (MMLU) zero-shot evaluation, perplexity on the Wikitext dataset (wiki), and perplexity on the Penn Treebank dataset (ptb). Our observations indicate significant variations in performance as a function of the parameter ratio and the compression method employed. Specifically, the table highlights the performance of each model when using standard SVD, SVD with binary search for truncation ranks (SVD*), and ASVD at different parameter ratios ranging from 0.75 to 0.95.
4.3 KV Cache Compression
Table 2: Performance under different KV cache compression ratio.
KV cache ratio
model dataset 1(original) 0.9 0.8 0.7 0.6 0.5 0.4 0.3 0.2
LLaMA-2-7b wiki 5.47 5.46 5.48 5.50 5.55 5.67 5.94 6.55 8.71
ptb 20.82 21.04 21.52 21.66 21.91 22.16 24.33 26.89 38.72
LLaMA-2-13b wiki 4.88 4.89 4.90 4.91 4.92 4.96 5.08 5.33 6.06
ptb 29.21 29.64 29.95 30.21 30.99 31.69 34.03 36.61 47.24
We evaluate the KV Cache compression by using ASVD to decompose k projection and v projection in transformer. Table 2 summarizes the results, showing the perplexities on the wikitext2 and Penn Treebank datasets for different KV cache compression ratios. It is evident that the perplexity values remain stable when the KV cache ratio is above 40%. When the ratio is lower than 40%, the performance of the network is decreased. These observations suggest that ASVD is effective to compress the KV cache without negatively impacting the model.
@@ 4.4 Integrating ASVD with Quantization
LLaMA-2-7b LLaMA-2-13b
param ratio dataset FP16 INT8 INT6 NF4 FP16 INT8 INT6 NF4
1 wiki 5.47 5.48 5.52 5.65 4.88 4.88 4.90 4.98
ptb 20.82 20.82 20.97 22.70 29.15 29.12 29.03 30.31
0.95 wiki 5.64 5.64 5.70 5.83 4.94 4.95 4.97 5.08
ptb 23.98 23.95 25.10 35.91 31.93 31.67 31.97 33.89
0.9 wiki 5.93 5.94 6.02 6.20 5.12 5.11 5.16 5.31
ptb 32.63 32.19 32.83 40.82 34.03 33.64 33.66 34.93
0.85 wiki 6.74 6.73 6.90 7.43 5.54 5.56 5.62 5.90
ptb 59.84 63.76 66.88 427.59 39.32 40.02 38.15 44.49
This section explores the integration of ASVD with quantization techniques to compress large language models. We investigate the synergy between ASVD and simple quantization methods, Round-To-Nearest (RTN) and 4-bit NormalFloat (NF4) [Dettmers et al., 2023], as a preliminary step. Future work could extend this investigation to more advanced quantization methods. Our experimental framework involves two stages. Firstly, we apply ASVD to decompose the network. Subsequently, we quantize the decomposed weights. We employ per-channel asymmetric quantization, targeting two quantization levels: 8-bit and 6-bit. We employ NF4 for quantizing to 4-bit.